Wentz, Inc. is considering going public but is unsure of a fair offering price for the...

  1. Wentz, Inc. is considering going public but is unsure of a fair offering price for the company. Before hiring an investment banker to assist in making the public offering, managers at Wentz have decided to make their own estimate of the firm’s common stock value. The firm’s CFO has gathered data for performing the valuating using the free cash flow valuation model.

      The firm’s weighted average cost of capital is 16%, and it has $2,000,000 of         debt at market value and $950,000 of preferred stock at its assumed market value. The estimated free cash flows over the next 5 year, 2013 through      2017, are given below. Beyond 2017 to infinity, the firm expects its free       cash flow to grow by 5% annually.

Year

FCF

2013

$125,000

2014

252,000

2015

314,525

2016

395,000

2017

455,750

  1. Estimate the value of Wentz Industries’ entire company by using the free cash flow valuation model.
  2. Use your finding in part a, along with the data provided above, to find Wentz Industries’ common stock value.
  3. If the firm plans to issue 125,000 shares of common stock, what is its estimated value per share?
Answer #1

1.

=125000/1.16+252000/1.16^2+314525/1.16^3+395000/1.16^4+455750/1.16^5+455750/1.16^5*1.05/(16%-5%)

=3002935.95

2.

=3002935.95-2000000-950000

=52935.95

3.

=52935.95/125000=0.4235